Protein Area (Your group) Area (Other group) Target 20318.945 48798.747 GAPDH 20157.359 20071.238 Ratio(Protein/GAPDH) ???? ????

Protein

Area (Your group)

Area (Other group)

Target

20318.945

48798.747

GAPDH

20157.359

20071.238

Ratio(Protein/GAPDH)

????

????
Answer #1

In this case, we can divide the Target by the GAPH. We will obtain the following

Ratio Area (your group) = 20318.945/20157.359 = 1.00

The ratio is 1:1 or 1/1

Ratio Area (other group) = 48798.747/20071.238 = 2.43

Approximately the ratio is 2:1 (2 Target: 1 GAPDH)